# Continuous optimization by quantum adaptive distribution search

### v1

@@ Abstract
In this paper, we introduce the quantum adaptive distribution search (QuADS), a quantum continuous optimization algorithm that integrates Grover adaptive search (GAS) with the covariance matrix adaptation - evolution strategy (CMA-ES), a classical technique for continuous optimization. QuADS utilizes the quantum-based search capabilities of GAS and enhances them with the principles of CMA-ES for more efficient optimization. It employs a multivariate normal distribution for the initial state of the quantum search and repeatedly updates it throughout the optimization process. Our numerical experiments show that QuADS outperforms both GAS and CMA-ES. This is achieved through adaptive refinement of the initial state distribution rather than consistently using a uniform state, resulting in fewer oracle calls. This study presents an important step toward exploiting the potential of quantum computing for continuous optimization.
@@ I INTRODUCTION
Inspired by the idea of adapting probability distributions in CMA-ES, we propose a natural extension of GAS, which we call quantum adaptive distribution search (QuADS), that combines Grover adaptive search and CMA-ES principles. Our algorithm aims to optimize continuous functions by adaptively changing the belief distribution and using the quantum-based search capabilities of amplitude amplification. Through numerical experiments in scenarios of up to 10 dimensions, we observe that GAS constantly needs high computational costs regardless of the type of function instead of the assurance of finding the global optimum. Also, CMA-ES typically converges faster than other methods, but it frequently falls into local optima. QuADS consistently outperforms GAS and performs better than CMA-ES, especially for high dimensional cases in terms of expected oracle counts until obtaining the global optimal solution. We believe QuADS will be one of the promising tools for accelerating continuous optimization on quantum computers.
The rest of this paper is organized as follows. In Section II, we provide summaries of GAS and CMA-ES. In Section III, we detail the main concepts and algorithms of QuADS. In Section IV, we validate our method through numerical experiments and show that the proposed method surpasses traditional approaches. Finally, we conclude the paper in Section V. Code for our numerical experiment can be found at Ref. [18].
@@ II.1 Grover Adaptive Search
GAS [6, 7] is a quantum optimization method that utilizes Grover’s algorithm [8, 9], which enables us to achieve quadratic speedup compared to classical random search. Notably, Grover’s algorithm is a specific instance of the more general amplitude amplification [19] (see Algorithm 1), with GAS corresponding to the case where the initial state is uniform. GAS begins with a uniform superposition state and, using the current minimum value as a threshold, amplifies the amplitudes of states below this threshold through amplitude amplification. See Algorithm 2 for the pseudo-code of the algorithm. $O_{f,\theta}$ is an oracle with objective function $f$ to perform a phase flip in regions where $f$ is less than the threshold $\theta$ and $H$ denotes Hadamard gate.
GAS has been extended to continuous optimization problems in Refs. [10, 11]. It is done by treating continuous optimization problems the same way as discrete ones by discretizing the continuous space using fixed-point representation. In continuous problems, the number of discretization grids is $2^{D\tau}$, where $D$ is the function’s dimension, and $\tau$ is the bit count for the continuous variable. This technique is especially effective for optimization problems of highly discontinuous functions due to the inherent nature of Grover’s algorithm, which accelerates the search in “needle in a haystack” scenarios.
@@
state preparation operator $P_{0}$, oracle $O_{f,\theta}$, threshold $\theta$
An obtained sample $x$, which satisfy $f(x)<\theta$
@@
       |$\ket{\psi}=\left(P_{0}\left(I-2\ket{0}\bra{0}\right)P_{0}^{\dagger}O_{f,\theta}\right)^{r}\ket{\psi}$
@@
until $f(x)<\theta$
@@
     $x\leftarrow\mathrm{AmplitudeAmplification}(H^{\otimes D},O_{f,\theta},\theta)$
     $y\leftarrow f(x)$
@@ II.2 CMA-ES
Adjust the step size to approximate the length of the evolved path to $\mathbb{E}\norm{\mathcal{N}(\bm{0},\bm{I})}$:
@@
Update $p_{c}$ to keep the average direction of $Z$:
@@
$C\leftarrow\left(1-c_{1}-c_{\mu}\sum_{j}w_{j}\right)C+c_{1}p_{c}p_{c}^{T}+c_{\mu}\sum_{i}^{K}w_{i}Z_{i}Z_{i}^{T}.$
@@
We briefly describe the CMA-ES algorithm (Algorithm 3). This method uses a multivariate normal distribution for the search process, denoted by $\mathcal{N}(\mu,\sigma^{2}C)$, where $\mu$ is the mean, $C$ is the scaled covariance, and $\sigma$ is the step size. While $C$ and $\sigma$ contribute to defining the covariance, $C$ primarily determines the shape and orientation, whereas $\sigma$ determines the scale of the search. The algorithm proceeds through the following iterative steps: Initially, $M$ search points, $x_{1},\ldots,x_{M}$, are generated from the normal distribution $\mathcal{N}(\mu,\sigma^{2}C)$. Subsequently, the function values $f(x_{1}),\ldots,f(x_{M})$ at each candidate point are evaluated, and the top $K(<M)$ individuals with the lowest function values are selected. Then, the parameters of the normal distribution are updated accordingly. $\mu$ is adjusted toward the average of the best individuals. $C$ is updated to reflect the direction in which the function value changes significantly, and $\sigma$ increases when $\mu$ consistently moves in a similar direction over several iterations and decreases when the movement of $\mu$ appears more random. For updating $C$, $\sigma$, the evolution path (moving average) $p_{c}$ and $p_{\sigma}$ are utilized. The detailed parameter update rule is shown in Algorithm 4 (see Ref. [14] for further detail). By repeating these operations, the normal distribution progressively approaches the optimal solution.
@@ III Quantum adaptive distribution search
The main concept of our method is to extend GAS by using an adaptive multivariate normal distribution for the initial state in quantum search. Denote $p$ is the probability of finding a point below a specified threshold in the distribution. Compared to GAS approach, QuADS is expected to achieve a higher probability $p$ by initiating the quantum search with an adaptive distribution instead of a uniform distribution. This is especially advantageous when the threshold is near the global optimum because while GAS consistently demands computational resources proportional to $O(\sqrt{2^{D\tau}})$, QuADS has the potential to reduce this cost when the adaptive distribution is properly updated toward the optima. From another perspective, our method enhances the efficiency of the sampling process in the CMA-ES by using amplitude amplification to sample from regions where function values are below an adaptive threshold. Classically, the number of oracle calls required to find a sample is $O(1/p)$. However, amplitude amplification can reduce this to $O(1/\sqrt{p})$ oracle calls. It should be noted that QuADS may occasionally converge on local optima due to its adaptive distribution updating strategy, in contrast to GAS, which always identifies the global optimum. Nonetheless, in terms of the expected oracle call to locate the global optimum, we experimentally find QuADS outperforms GAS, as will be discussed in detail in Section IV.
@@
where $Z$ is the normalization constant. The construction of $\mathcal{G}_{\mu,\Sigma}$ has been discussed in Refs. [21, 22]. This operator can be implemented using a polynomial number of quantum gates with respect to $D$ and $\tau$. Our approach begins by sampling $K$ superior points from the region below the adaptive threshold $\theta$. In the sampling, amplitude amplification is used, with the normal distribution serving as an initial state. Figure 1 illustrates the corresponding quantum circuit for the amplitude amplification. These sampled points are then used to update both the thresholds and the normal distribution parameters. For updating the distribution parameters, we employ the same updating rule as in CMA-ES, owing to the invariance properties that are also applicable in our framework. By utilizing the updating rule from CMA-ES, we can combine insights about the structure of the function and lead to an efficient search. As for the updating rule for thresholds, QuADS uses a moving average of the $q$-quantile of the function value at the sampled points, i.e., approximately equivalent to considering the best $Mq$th value out of $M$ samples, for some constant $q$. The reason for using this threshold update rule rather than using the minimum of sampled values as done in GAS is that if the threshold is updated prematurely, the distribution may not have sufficient weights on the region below the threshold, and that causes inefficient sampling. Fig. 2 demonstrates a typical optimization process using QuADS.
@@
Figure 2: Typical optimization process of QuADS. The solid line represents the $1\sigma$ region of the distribution at each iteration, and the points of the same color as a line represent the samples from amplitude amplification using the distribution as the initial state. The color becomes darker as iteration progresses.
IV Experiments
@@ IV.1 Performance estimation by amplitude simulation
For multiple test functions shown in Table 1, we run each method for $100$ times. We scaled the domain of the $D$ dimensional test function to be $[0,1]^{D}$. In CMA-ES and QuADS, we uniformly sampled the initial mean $\mu_{0}$ from the function domain for each trial. We initialized the covariance matrix, $C_{0}$, as the identity matrix, and set the initial step size, $\sigma_{0}$, to 0.5. We used $f(\mu_{0})$ as the initial threshold of QuADS. For the threshold of GAS, we used $f(x)$ as the initial threshold, where $x$ is a sample from a uniform distribution on the domain. The termination conditions for CMA-ES and QuADS are local convergence or finding the global optimum. Finding the global optimum is defined as sampling a point in the $\epsilon=0.01$ neighborhood of the global optimum. Note that the global optimum of the test functions is known. Also, we regard the local convergence as achieved when $\sigma$ becomes smaller than $\epsilon_{\sigma}=0.01$. For updating normal distribution in QuADS and CMAES, we use the default hyper-parameters in Ref. [14]. For the threshold update in QuADS, we heuristically use $\alpha=0.5$ and $q=0.2$. For the number of samples $M$ and the number of selected samples $K$ in CMA-ES, we use the default setting in Ref. [14],
@@
For QuADS, we use the same $K$ as CMA-ES. Finally, for amplitude amplification, we use an increase rate of $\lambda=6/5$.
@@
Name Function form
rastrigin $\sum_{i=1}^{D}\left(x_{i}^{2}-10\cos\left(2\pi x_{i}\right)\right)$
ackley $-20\exp\left(-0.2\sqrt{\frac{1}{D}\sum_{i=1}^{D}x_{i}^{2}}\right)-\exp\left(\frac{1}{D}\sum_{i=1}^{D}\cos\left(2\pi x\right)\right)$
styblinski tang $\frac{1}{2}\sum_{i=1}^{D}\left(x_{i}^{4}-16x_{i}^{2}+5x_{i}\right)$
schwefel $\sum_{i=1}^{D}\left(x_{i}\sin\sqrt{\absolutevalue{x_{i}}}\right)$
griewank $\frac{1}{4000}\sum_{i=1}^{n}x_{i}^{2}-\prod_{i=1}^{n}\cos\left(\sqrt{x_{i}}\right)$
rosenbrock $\sum_{i=1}^{D-1}\left(100\left(x_{i+1}-x_{i}^{2}\right)^{2}+\left(1-x_{i}\right)^{2}\right)$
alpine01 $\sum_{i=1}^{D}\absolutevalue{x_{i}\sin\left(x_{i}\right)+\frac{1}{10}x_{i}}$ |
alpine02 $-\prod_{i=1}^{D}\left(\sqrt{x_{i}}\sin\left(x_{i}\right)\right)$
deflected corrugated spring $\frac{1}{10}\sum_{i=1}^{D}\left(x_{i}^{2}-\cos\left(5\sqrt{\sum_{j=1}^{D}x_{j}^{2}}\right)\right)$
wavy $-\sum_{i=1}^{D}\cos\left(10x_{i}\right)\exp\left(-x_{i}^{2}/2\right)$
@@
where $o_{\rm local}$ and $o_{\rm global}$ are the average number of oracle calls used to reach local or global convergences, respectively, and $p_{\rm global}$ is the probability to find the global optimum. Note that we can approximate it as $p_{\rm global}=n_{\rm global}/100$ where $n_{\rm global}$ is the number of experiments in which we found the global optimum. Then, the expected oracle calls until global convergence can be calculated as
@@
Fig. 3 shows $o_{\rm total}$ of each method. The data clearly show that QuADS consistently outperforms GAS in all cases. Compared to CMA-ES, our approach demonstrates at least comparable performance and, notably, it significantly outperforms CMA-ES on functions that are challenging for CMA-ES the presence of numerous similar local optima in the neighborhood of the global optimum.
@@
Figure 3: Expected oracle call counts for finding global optimum for each method (Eq. (2)) when $D=3$. We utilized the bootstrap re-sampling method to generate the error bars in these figures. This method involves randomly selecting data from the pool of 100 experiments and recalculating Eq. (2) for each of these selected data sets. We present the estimated 5th and 95th percentiles of the expected oracle evaluation count as confidence intervals. Confidence intervals outside the figure indicate that upper bounds cannot be estimated from bootstrap.
@@
Figure 4: Optimization processes for the three-dimensional rastrigin function (4a) and schwefel function (4b). The upper figure presents the smallest function value obtained up to a specific number of oracle calls in each trial. A circle marker at the end of each trial represents the trial found a global solution, and a cross marker indicates local convergence. The lower figure shows the proportion of trials that found a global solution (solid line) and that reached the terminal condition (dashed line) up to the number of oracle calls.

### v2

@@ Abstract
In this paper, we introduce the quantum adaptive distribution search (QuADS), a quantum continuous optimization algorithm that integrates Grover adaptive search (GAS) with the covariance matrix adaptation - evolution strategy (CMA-ES), a classical technique for continuous optimization. QuADS utilizes the quantum-based search capabilities of GAS and enhances them with the principles of CMA-ES for more efficient optimization. It employs a multivariate normal distribution for the initial state of the quantum search and repeatedly updates it throughout the optimization process. Our numerical simulations show that QuADS outperforms both GAS and CMA-ES. This is achieved through adaptive refinement of the initial state distribution rather than consistently using a uniform state, resulting in fewer oracle calls. This study presents an important step toward exploiting the potential of quantum computing for continuous optimization.
@@ I INTRODUCTION
Inspired by the idea of adapting probability distributions in CMA-ES, we propose a natural extension of GAS, which we call quantum adaptive distribution search (QuADS), that combines Grover adaptive search and CMA-ES principles. Our algorithm aims to optimize continuous functions by adaptively changing the belief distribution and using the quantum-based search capabilities of amplitude amplification. Through numerical simulations in scenarios of up to 10 dimensions, we observe that GAS constantly needs high computational costs regardless of the type of function instead of the assurance of finding the global optimum. Also, CMA-ES typically converges faster than other methods, but it frequently falls into local optima. QuADS consistently outperforms GAS and performs better than CMA-ES, especially for high dimensional cases in terms of expected oracle counts until obtaining the global optimal solution. We believe QuADS will be one of the promising tools for accelerating continuous optimization on quantum computers.
The rest of this paper is organized as follows. In Section II, we provide summaries of GAS and CMA-ES. In Section III, we detail the main concepts and algorithms of QuADS. In Section IV, we validate our method through numerical simulations and show that the proposed method surpasses traditional approaches. Finally, we conclude the paper in Section V. Code for our numerical simulation can be found at Ref. [18].
@@ II.1 Grover Adaptive Search
Grover’s algorithm [8, 9] achieves a quadratic speedup in searching for specific items within unstructured tables. Specifically, given a set of elements $A=\{a_{i}\}$ $(i=0,\dots,n-1)$ and a subset of ’good’ elements $G\subset A$, Grover’s algorithm efficiently identifies the index of the element that belong to $G$. Grover’s algorithm is a special case of the more general concept of amplitude amplification [19], where the initial state is uniform. The algorithm for amplitude amplification is shown in Algorithm 1. $O$ represents an oracle gate, which selectively inverts the phase of the amplitude of states corresponding to the ’good’ elements. $\lambda$ is selected in the range $(1,4/3)$ because it is guaranteed to achieve quadratic speedup within this range [9].
Grover Adaptive Search (GAS) [6, 7] is a quantum optimization method that utilizes Grover’s algorithm. GAS begins with a uniform superposition state and, using the current minimum value as a threshold, amplifies the amplitudes of states below this threshold through Grover’s search. See Algorithm 2 for the pseudo-code of the algorithm. $O_{f,\theta}$ is an oracle with objective function $f$ to perform a phase flip in regions where $f$ is less than the threshold $\theta$ and $H$ denotes Hadamard gate.
GAS has been extended to continuous optimization problems in Refs. [10, 11]. It is done by treating continuous optimization problems the same way as discrete ones by discretizing the continuous space using fixed-point representation. In continuous problems, the number of discretization grids is $2^{D\tau}$, where $D$ is the function’s dimension, and $\tau$ is the bit count for the continuous variable. This technique is especially effective for optimization problems of highly discontinuous functions due to the inherent nature of Grover’s algorithm, which accelerates the search in “needle in a haystack” scenarios. However, in practical situations, GAS is not the best choice because employing methods beyond uniform search is advantageous.
@@
state preparation operator $P_{0}$, oracle $O$
@@
[…]
@@
until good element is found
@@
     $x\leftarrow\mathrm{AmplitudeAmplification}(H^{\otimes D},O_{f,\theta})$
     $\theta\leftarrow f(x)$
@@ II.2 CMA-ES
Adjust the step size to keep the length of the evolved path close to $\mathbb{E}\norm{\mathcal{N}(\bm{0},\bm{I})}$:
@@
Define $h_{\sigma}^{g}$ where $g$ denote current iteration number. $h_{\sigma}^{g}$ is used to prevent $C$ from increasing too fast.
$h_{\sigma}:=\begin{cases}1&\text{if }\frac{\norm{p_{\sigma}}}{\sqrt{1-(1-c_{\sigma})^{2(g+1)}}}<(1.4+\frac{2}{n+1})\mathbb{E}\norm{N(0,1)}\\ 0&\text{otherwise}\end{cases}$
Update $p_{c}$ to keep the average direction of $Z$.
@@
$\begin{split}C\leftarrow&\left(1+c_{1}\delta(h_{\sigma})-c_{1}-c_{\mu}\sum_{j=1}^{K}w_{j}\right)C\\ &+c_{1}p_{c}p_{c}^{T}+c_{\mu}\sum_{i=1}^{K}w_{i}Z_{i}Z_{i}^{T}.\end{split}$
@@
We briefly describe the CMA-ES algorithm (Algorithm 3). This method uses a multivariate normal distribution for the search process, denoted by $\mathcal{N}(\mu,\sigma^{2}C)$, where $\mu$ is the mean, $C$ is the scaled covariance, and $\sigma$ is the step size. While $C$ and $\sigma$ contribute to defining the covariance, $C$ primarily determines the shape and orientation, whereas $\sigma$ determines the scale of the search. The algorithm proceeds through the following iterative steps: Initially, $M$ search points, $x_{1},\ldots,x_{M}$, are generated from the normal distribution $\mathcal{N}(\mu,\sigma^{2}C)$. Subsequently, the function values $f(x_{1}),\ldots,f(x_{M})$ at each candidate point are evaluated, and the top $K(<M)$ individuals with the lowest function values are selected. Then, the parameters of the normal distribution are updated accordingly. $\mu$ is adjusted toward the average of the best individuals. $C$ is updated to reflect the direction in which the function value changes significantly, and $\sigma$ increases when $\mu$ consistently moves in a similar direction over several iterations and decreases when the movement of $\mu$ appears more random. For updating $C$ and $\sigma$, the evolution path (moving average) $p_{c}$ and $p_{\sigma}$ are utilized. The detailed parameter update rule is shown in Algorithm 4 (see Ref. [14] for further detail). By repeating these operations, the normal distribution progressively approaches the optimal solution.
@@ III Quantum adaptive distribution search
The main concept of our method is to extend GAS by using an adaptive multivariate normal distribution for the initial state in quantum search. Denote $p$ is the probability of finding a point below a specified threshold in the distribution. Compared to GAS approach, QuADS is expected to achieve a higher $p$ by initiating the quantum search with an adaptive distribution instead of a uniform distribution. This is especially advantageous when the threshold is near the global optimum because while GAS consistently demands computational resources proportional to $O(\sqrt{2^{D\tau}})$, QuADS has the potential to reduce this cost when the adaptive distribution is properly updated toward the optima. From another perspective, our method enhances the efficiency of the sampling process in the CMA-ES by using amplitude amplification to sample from regions where function values are below an adaptive threshold. Classically, the number of oracle calls required to find a sample is $O(1/p)$. However, amplitude amplification can reduce this to $O(1/\sqrt{p})$ oracle calls. It should be noted that QuADS may occasionally converge on local optima due to its adaptive distribution updating strategy, in contrast to GAS, which always identifies the global optimum. Nonetheless, in terms of the expected oracle call to locate the global optimum, we find QuADS outperforms GAS, as will be discussed in detail in Section IV.
@@
where $Z$ is the normalization constant. The construction of $\mathcal{G}_{\mu,\Sigma}$ has been discussed in Refs. [21, 22]. This operator can be implemented using a polynomial number of quantum gates with respect to $D$ and $\tau$. Our approach begins by sampling $K$ superior points from the region below the adaptive threshold $\theta$. In the sampling, amplitude amplification is used, with the normal distribution serving as an initial state. Figure 1 illustrates the corresponding quantum circuit for the amplitude amplification. These sampled points are then used to update both the thresholds and the normal distribution parameters. For updating the distribution parameters, we employ the same updating rule as in CMA-ES. By utilizing the updating rule from CMA-ES, we can combine insights about the structure of the function and lead to an efficient search. As for the updating rule for thresholds, QuADS uses a moving average of the $q$-quantile of the function value at the sampled points, i.e., approximately equivalent to considering the best $Mq$th value out of $M$ samples, for some constant $q$. The reason for using this threshold update rule rather than using the minimum of sampled values as done in GAS is that if the threshold is updated prematurely, the distribution may not have sufficient weights on the region below the threshold, and that causes inefficient sampling.
Fig. 2 shows an example of the QuADS optimization process for the alpine02 function in Table 1. The mean of the normal distribution progressively shifts toward the global optimum, while the distribution’s covariance broadens in the direction of this movement. Each iteration’s points are sampled through amplitude amplification using the normal distribution as a prior. Notably, more samples are drawn from areas where the normal distribution’s weights are larger, within regions with small function values.
@@
Figure 2: Typical optimization process of QuADS. The solid line represents the $1\sigma$ region of the distribution, and the points represent the samples from amplitude amplification at each iteration. The star mark represents the global optimum point.
IV Numerical Simulations
@@ IV.1 Performance estimation by amplitude simulation
For multiple test functions shown in Table 1, we run each method for $100$ times. We chose these test functions for their multimodal characteristics and their ability to be defined in arbitrary dimensions. We scaled the domain of the $D$ dimensional test function to be $[0,1]^{D}$. The plots of the test functions are shown in Appendix C. In CMA-ES and QuADS, we uniformly sampled the initial mean $\mu_{0}$ from the function domain for each trial. We initialized the covariance matrix, $C_{0}$, as the identity matrix, and set the initial step size, $\sigma_{0}$, to 0.5. We used $f(\mu_{0})$ as the initial threshold of QuADS. For the threshold of GAS, we used $f(x)$ as the initial threshold, where $x$ is a sample from a uniform distribution on the domain. The termination conditions for CMA-ES and QuADS are local convergence or finding the global optimum. Finding the global optimum is defined as sampling a point in the $\epsilon=0.01$ neighborhood of the global optimum. Note that the global optimum of the test functions is known. Also, we regard the local convergence as achieved when $\sigma$ becomes smaller than $\epsilon_{\sigma}=0.01$. For updating normal distribution in QuADS and CMA-ES, we use the hyperparameters shown in Appendix A. For the threshold update in QuADS, we heuristically use $\alpha=0.5$ and $q=0.2$. For the number of samples $M$ and the number of selected samples $K$ in CMA-ES, we use the default setting in Ref. [14],
@@
For QuADS, we use the same $K$ as CMA-ES. Finally, for amplitude amplification, we set the increase rate $\lambda$ to $6/5$, suggested in [9].
@@
Name Function form Domain
rastrigin $\sum_{i=1}^{D}\left(x_{i}^{2}-10\cos\left(2\pi x_{i}\right)\right)$ [-5.12, 5.12]
ackley $-20\exp\left(-0.2\sqrt{\frac{1}{D}\sum_{i=1}^{D}x_{i}^{2}}\right)-\exp\left(\frac{1}{D}\sum_{i=1}^{D}\cos\left(2\pi x\right)\right)$ [-4, 4]
styblinski tang $\frac{1}{2}\sum_{i=1}^{D}\left(x_{i}^{4}-16x_{i}^{2}+5x_{i}\right)$ [-5, 5]
schwefel $\sum_{i=1}^{D}\left(x_{i}\sin\sqrt{\absolutevalue{x_{i}}}\right)$ [-500, 500]
griewank $\frac{1}{4000}\sum_{i=1}^{n}x_{i}^{2}-\prod_{i=1}^{n}\cos\left(\sqrt{x_{i}}\right)$ [-512, 512]
alpine01 $\sum_{i=1}^{D}\absolutevalue{x_{i}\sin\left(x_{i}\right)+\frac{1}{10}x_{i}}$ | [-10, 10]
alpine02 $-\prod_{i=1}^{D}\left(\sqrt{x_{i}}\sin\left(x_{i}\right)\right)$ [0, 10]
deflected corrugated spring $\frac{1}{10}\sum_{i=1}^{D}\left(x_{i}^{2}-\cos\left(5\sqrt{\sum_{j=1}^{D}x_{j}^{2}}\right)\right)$ [0, 10]
wavy $-\sum_{i=1}^{D}\cos\left(10x_{i}\right)\exp\left(-x_{i}^{2}/2\right)$ [-$\pi$, $\pi$]
@@
where $o_{\rm local}$ and $o_{\rm global}$ are the average number of oracle calls used to reach local or global convergences, respectively, and $p_{\rm global}$ is the probability to find the global optimum. Note that we can approximate it as $p_{\rm global}=n_{\rm global}/100$ where $n_{\rm global}$ is the number of simulations in which we found the global optimum. Then, the expected oracle calls until global convergence can be calculated as
@@
Fig. 3 shows $o_{\rm total}$ of each method. Additionally, the figure includes results for two well-established classical optimizers: particle swarm optimization (PSO) [23, 24] and basinhopping optimizer [25] for reference. The hyperparameters and implementations for the PSO and basinhopping optimizer are detailed in Appendix B. The data clearly show that QuADS consistently outperforms GAS in all cases. For all functions except the griewank function, GAS shows almost the same number of oracle call counts. The reason for the increased oracle call count in the griewank function is due to its significant fluctuations around the global optimum (See Appendix C), resulting in a smaller success region, where the distance from the global optimum is within $\epsilon_{\sigma}$ and the function value is below the threshold. Compared to classical methods CMA-ES, PSO and basinhopping, our approach demonstrates at least comparable performance and, notably, it surpasses them on rastrigin, schwefel, styblinski tang, alpine02, wavy function. These functions, other than the rastrigin function, are characterized by multiple big valleys within the search space, each containing promising local solutions, whose optimization is difficult for CMA-ES. On the other hand, in the case of the ackley, alpine01, and griewank functions, characterized by numerous small local optima within a large valley, QuADS demonstrates performance nearly identical to that of CMA-ES in a three-dimensional setup. This structure makes optimization more straightforward for CMA-ES, leading to an increased $p$, which consequently reduces the influence of quadratic acceleration. However, in scenarios with higher dimensions, where $p$ is likely to be lower, the benefits of quantum search would become more pronounced.
@@
Figure 3: Expected oracle call counts for finding global optimum for each method (Eq. (2)) when $D=3$. We utilized the bootstrap re-sampling method to generate the error bars in these figures. This method involves randomly selecting data from the pool of 100 simulations and recalculating Eq. (2) for each of these selected data sets. We present the estimated 5th and 95th percentiles of the expected oracle evaluation count as confidence intervals. Confidence intervals outside the figure indicate that upper bounds cannot be estimated from bootstrap.
@@
Figure 4: Optimization processes for the three-dimensional rastrigin function (4a) and schwefel function (4b). The upper figure presents the smallest function value obtained up to a specific number of oracle calls in each trial. A circle marker at the end of each trial represents the trial found a global solution, and a cross marker indicates local convergence. The lower figure shows the proportion of trials that found a global solution (solid line) and that reached the terminal condition (dashed line) up to the number of oracle calls. The number of oracle calls required for local convergence can be larger than those for finding the global optimum. This arises from the algorithms’ termination criteria: they terminate immediately upon locating the global optimum, but may continue until convergence of $\sigma$ if the global optimum remains unfound.